OpenStudy (anonymous):

solve the next separable ODE and graph the direction field of all solutions...

OpenStudy (anonymous):



OpenStudy (anonymous):

Starting with \[\frac{dy}{dx}=\frac{1-x^2}{y^2}\] Multiply both sides by y^2 \[y^2\frac{dy}{dx}=1-x^2\] Multiply both sides by \[y^2dy=(1-x^2)dx\] Integrate both sides \[\int\limits y^2dy=\int\limits (1-x)dx\]Find the integral and solve for y.

OpenStudy (anonymous):

i have troubles in the direction field \[y = \sqrt(1-x^2+c)\]

OpenStudy (amistre64):

the direction field is defined by dy/dx

OpenStudy (amistre64):

come up with some values for x and y and plug them into dy/dx to define the slope at the point

OpenStudy (anonymous):

with initial conditions?

OpenStudy (anonymous):

could you draw it?

OpenStudy (amistre64):

i cant seem to do much of anything today ... sites gone wacky

OpenStudy (amistre64):

say we want a 10x10 grid such that -5<=x<=5; and same for y -5 -4 -3 -2 -1 0 1 2 3 4 5 -5 (-5,-5) (-5,-4) (-5,-3) .... -4 (-4,-5) (-4,-4) (-4,-3) .... -3 ... -2 -1 0 1 2 3 4 5 the poitns themselves define the x and y values to plug into dy/dx to determine the slopes with

OpenStudy (amistre64):

\[\frac{dy}{dx}=\frac{1-(-5)^2}{(-5)^2}=-24/25\]almost a slope of -1 at the point -5,-5

OpenStudy (amistre64):

at x=-1 or 1, the slope = 0

OpenStudy (anonymous):

i am studying the method of the Isoclines, anyway thanks

OpenStudy (amistre64):